# PosIR: Position-Aware Heterogeneous Information Retrieval Benchmark

### v1

@@ 6 Conclusion
In this paper, we introduce a new IR benchmark, PosIR, designed to diagnose position bias in information retrieval models. Unlike existing evaluation suites, PosIR leverages granular reference spans and length-diverse settings to effectively disentangle the effects of document length from the position of relevant information. Our experiments expose prevalent primacy and recency biases across both multilingual and cross-lingual retrieval settings, underscoring the limitations of current evaluation practices. By providing a reproducible pipeline and a diverse suite of 310 datasets, PosIR fills a crucial gap in the evaluation ecosystem, paving the way for more position-robust retrieval systems.
@@
Despite our efforts, this study has several limitations: 1) Although we employ LLMs to extract reference spans and implement a strict contrastive verification pipeline, a marginal risk remains that certain alternative reference spans might be overlooked. While our quality control is designed for high precision, such omissions could introduce noise into the position-aware evaluation. 2) While LLM-based translation significantly extends the linguistic diversity of PosIR, it is intrinsically difficult to completely eliminate subtle semantic drift or translation artifacts across all datasets. Although our sampling-based automatic and human evaluations indicate high fidelity, translation quality in certain low-resource languages might still be influenced by the underlying model bias. 3) Our empirical analysis focuses on embedding-based dense retrieval models. Although prior research Zeng et al. (2025b) indicates that cross-encoders may exhibit superior position robustness, a systematic large-scale investigation of alternative architectures, including generative retrieval, remains for future work. Nevertheless, PosIR provides the necessary model-agnostic framework to facilitate such benchmarking. 4) Our investigation is currently limited to text retrieval. Whether and how position bias manifests in multimodal retrieval contexts (e.g., text-to-image or video retrieval) remains an open question. We leave the exploration of these cross-modal positional dynamics to future work.
Ethical Considerations
The construction of PosIR involves text generation by LLMs. Consequently, the synthetic components of the datasets (e.g., queries) may inherently reflect the socio-cultural biases, stereotypes, or toxicity present in the base models. Additionally, as the document corpora are derived from real-world open-source datasets (IndustryCorpus2 and FineWeb), they may inevitably contain personally identifiable information or sensitive content. To mitigate these risks, we emphasize that PosIR is intended solely for research benchmarking and evaluation purposes. We strictly prohibit its use for training generative models to prevent the propagation of potential toxicity or hallucinations.
Appendix A PosIR Datasets
A.1 Overview
@@
PosIR comprises 310 datasets (17,329,673 documents and 421,708 queries) spanning 10 languages and 31 domains, forming a large-scale multilingual benchmark for information retrieval. For each dataset, we use the same format as BEIR, i.e. corpus, queries and qrels, which are all available in the Hugging Face Hub. [^15] https://huggingface.co/datasets/infgrad/PosIR-Benchmark-v1 Table 4 summarizes language-level statistics, showing well-balanced corpus sizes across languages, with each language contributing approximately 1.7M–1.9M documents. Token length statistics reveal substantial cross-lingual variation that reflects differences in writing systems and morphological complexity. Logographic languages exhibit more compact representations (e.g., Chinese and English), whereas morphologically rich alphabetic languages require longer sequences (e.g., Arabic and Russian). Table 5 reports domain-level aggregates and highlights pronounced variation in domain scale and document characteristics. Domain sizes range from several hundred thousand documents to over three quarters of a million, while token length distributions reflect domain-specific content properties. Technical domains such as Information Security tend to contain longer documents, whereas more formal or symbolic domains, such as Mathematics and Statistics, exhibit comparatively shorter texts. This diversity enables evaluation of IR systems under heterogeneous conditions. For comprehensive coverage, Tables 1343 provide detailed statistics for all 31 domains, presented individually for each domain. These tables report language-specific document counts, query counts, average token lengths, and token length ratios relative to English. Such fine-grained reporting supports reproducibility and facilitates the analysis of domain–language interaction effects in multilingual IR.
@@
Table 5: Summary of corpus statistics aggregated by domain.
A.2 Diversity Analysis
Query Diversity
@@ A.2 Diversity Analysis
Corpus Diversity
Following the approach in MTEB Muennighoff et al. (2023), we compute inter-domain similarities separately for English and Chinese corpora, each consisting of 31 domains, in PosIR. Specifically, we adopt the highest-scoring model on the MTEB (eng, v2) STS task, QZhou-Embedding (91.65) Yu et al. (2025), to generate embeddings for all documents within each domain. We then average document embeddings at the domain level and compute pairwise cosine similarities between domains. The resulting similarity matrix is visualized as a heatmap in Figure 6.
@@
Appendix B Multilingual Machine Translation
@@
We evaluate four representative translation systems: (1) Google Translate [^17] https://translate.google.com, a widely used commercial neural machine translation service; (2) Hunyuan-MT-7B Zheng et al. (2025), a specialized translation model achieved first place in the WMT25 competition; (3) Qwen3-30B-A3B-Instruct-2507 Yang et al. (2025a), a 31B-parameter open-source general MoE model; and (4) GPT-4o OpenAI et al. (2024), a frontier general model serving as a strong baseline. Each system translates the sampled texts into 9 target languages: Arabic (ara-Arab), French (fra-Latn), German (deu-Latn), Italian (ita-Latn), Korean (kor-Kore), Portuguese (por-Latn), Russian (rus-Cyrl), and Spanish (spa-Latn). The prompt for the translation task is provided in Appendix C.4.
B.1 Automatic Evaluation
B.1.1 Evaluation Protocol
To efficiently evaluate the translation quality across 8 languages and 31 domains, we employ an LLM-based reference-free evaluation approach Kocmi and Federmann (2023) (see Appendix C.5). Specifically, we utilize GPT-4o as an automatic evaluator to assess translation quality along four dimensions: Fluency (naturalness and readability of the target text), Accuracy (semantic fidelity to the source), Completeness (preservation of all source information), and Style Consistency (appropriateness of tone and register). Each dimension is rated on a 5-point Likert scale (1 = Poor, 5 = Excellent), and an overall score is computed as the mean of the four dimensions. In total, this results in 2,976 evaluated translation outputs (372 samples $\times$ 8 languages).
@@ B.1.2 Evaluation Results
B.2 Human Evaluation
Initial inspection using automatic evaluation suggests that Qwen3-30B-A3B-Instruct-2507 produces high-quality translations across PosIR, achieving competitive performance in comparison to other methods. However, given that translation quality appears strong, rigorous human validation is essential. While automatic evaluation can detect catastrophic failures, it may be insensitive to subtle but critical errors in high-quality translations, such as nuanced mistranslations, contextual inappropriateness, or domain-specific inaccuracies (Freitag et al., 2021). Due to limited resources, we conduct focused human evaluation on three typologically diverse languages (Arabic, French, and Russian) to provide expert assessment of translation quality, facilitating a direct comparison between human evaluation and LLM-based assessments.
@@ B.2 Human Evaluation
B.2.1 Annotator Recruitment
Five annotators were recruited from graduate programs in linguistics, natural language processing, and applied foreign languages via established academic networks. All annotators were advanced graduate students or recent PhD graduates with expertise in computational linguistics and/or translation studies. For Arabic and Russian, native speakers of the respective languages conducted the annotations. For French, two annotators holding DALF C1/C2 certificates (Diplôme Approfondi de Langue Française, the highest level of French language certification) completed the evaluation. All annotators were compensated at the French SMIC (Salaire Minimum Interprofessionnel de Croissance) hourly rate of €12.02 (\$14.15 USD as of December 2025), in compliance with French labor regulations.
B.2.2 Evaluation Protocol
@@
Annotators evaluated 372 translation pairs across 3 languages (1,116 samples in total) generated by Qwen3-30B-A3B-Instruct-2507. The evaluations were conducted in accordance with the consistency standards outlined in the automatic evaluation, as described in Appendix B.1.1. Annotators were given both the English source text and the target language translation, with no access to reference translations to prevent anchoring bias.
B.2.3 Evaluation Results
@@
Table 9: Human-GPT-4o agreement within ±1 point tolerance.
B.3 GPT-4o Human Evaluation Alignment
@@
B.4 Model Selection Rationale
@@
B.5 Fixing and Filtering Strategies
@@ B.5.2 Refusal Detection
B.5.3 Statistical Outlier Filtering
We perform token-length-based outlier filtering at the language–domain level using the Qwen3 tokenizer. For each stratum, we compute the interquartile range (IQR) and define asymmetric thresholds, with a lower bound of $Q1-3.0\times IQR$ and an upper bound of $Q3+1.5\times IQR$. This asymmetric design focuses on filtering abnormally long texts, which are more likely to result from translation errors, duplication, or unintended document concatenation. It enforces a stricter upper bound to remove such cases, while allowing reasonable length variation among shorter documents. We process only corpus outliers, cascading deletions to associated queries and qrels. In total, 15,641 corpus documents (0.11%) and 332 queries (0.10%) are removed. The largest reductions occur in Arabic (5,490 corpus entries; 0.25%) and Korean (4,082; 0.24%), followed by Russian (1,557; 0.11%). Overall, the removal ratios remain consistently low across all languages, indicating that the filtering procedure effectively eliminates anomalous long texts while preserving the structural integrity and cross-lingual balance of the benchmark.
Appendix C Prompts
C.1 Prompt for Position-Aware QA Generation
@@
C.2 Prompt for Reference Locating
@@
C.3 Prompt for Relevance Assessment
Thomas et al. (2024) demonstrate that large language models such as GPT-4 can achieve labeling quality comparable to human annotators when producing gold-standard relevance labels for search systems. Moreover, Zhuang et al. (2024); Zeng et al. (2025a) show that incorporating fine-grained relevance labels into prompts for LLM-based rerankers significantly improves zero-shot reranking performance by better distinguishing borderline and ambiguous relevance cases. Motivated by these findings, we employ DeepSeek-V3.1 (Non-thinking Mode) [^20] https://api-docs.deepseek.com/news/news250821 as a relevance labeler, adopting a 5-level relevance labeling scheme to capture fine-grained distinctions in relevance. This labeling process is used solely for auxiliary verification rather than primary relevance annotation.
@@ C.3 Prompt for Relevance Assessment
C.4 Prompt for Multilingual Machine Translation
@@
C.5 Prompt for Translation Quality Evaluation
@@
Qwen3 tokenizer Yang et al. (2025a) - https://huggingface.co/Qwen/Qwen3-30B-A3B-Instruct-2507/blob/main/tokenizer.json
@@
gte-multilingual-base Zhang et al. (2024) 305M https://huggingface.co/Alibaba-NLP/gte-multilingual-base
bge-m3 Chen et al. (2024) 568M https://huggingface.co/BAAI/bge-m3
Qwen3-Embedding-0.6B Zhang et al. (2025) 595M https://huggingface.co/Qwen/Qwen3-Embedding-0.6B
inf-retriever-v1-1.5b Yang et al. (2025b) 2B https://huggingface.co/infly/inf-retriever-v1-1.5b
Qwen3-Embedding-4B Zhang et al. (2025) 4B https://huggingface.co/Qwen/Qwen3-Embedding-4B
inf-retriever-v1 Yang et al. (2025b) 7B https://huggingface.co/infly/inf-retriever-v1
QZhou-Embedding Yu et al. (2025) 7B https://huggingface.co/Kingsoft-LLM/QZhou-Embedding
NV-Embed-v2 Lee et al. (2025) 8B https://huggingface.co/nvidia/NV-Embed-v2
llama-embed-nemotron-8b Babakhin et al. (2025) 8B https://huggingface.co/nvidia/llama-embed-nemotron-8b
Qwen3-Embedding-8B Zhang et al. (2025) 8B https://huggingface.co/Qwen/Qwen3-Embedding-8B
KaLM-Embedding-Gemma3-12B-2511 Zhao et al. (2025a) 12B https://huggingface.co/tencent/KaLM-Embedding-Gemma3-12B-2511
@@
bge-reranker-v2-m3 Chen et al. (2024) 0.6B https://huggingface.co/BAAI/bge-reranker-v2-m3
Qwen3-Reranker-0.6B Zhang et al. (2025) 0.6B https://huggingface.co/Qwen/Qwen3-Reranker-0.6B
bge-reranker-v2-minicpm-layerwise Chen et al. (2024) 3B https://huggingface.co/BAAI/bge-reranker-v2-minicpm-layerwise
Qwen3-Reranker-4B Zhang et al. (2025) 4B https://huggingface.co/Qwen/Qwen3-Reranker-4B
@@
Hunyuan-MT-7B Zheng et al. (2025) 8B https://huggingface.co/tencent/Hunyuan-MT-7B
Qwen3-30B-A3B-Instruct-2507 Yang et al. (2025a) 31B https://huggingface.co/Qwen/Qwen3-30B-A3B-Instruct-2507
DeepSeek-V3.1 DeepSeek-AI et al. (2025) 685B https://huggingface.co/deepseek-ai/DeepSeek-V3.1
GPT-4o OpenAI et al. (2024) - https://platform.openai.com/docs/models/gpt-4o
@@
Table 17: Statistics for domain: Automobile (Part 5 of 31).
@@
Table 22: Statistics for domain: Electric Power Energy (Part 10 of 31).
@@
Table 23: Statistics for domain: Film Entertainment (Part 11 of 31).
@@
Table 24: Statistics for domain: Finance Economics (Part 12 of 31).
@@
Table 27: Statistics for domain: Game (Part 15 of 31).
@@
Table 28: Statistics for domain: Law Judiciary (Part 16 of 31).
@@
Table 29: Statistics for domain: Literature Emotion (Part 17 of 31).
@@
Table 32: Statistics for domain: Mining (Part 20 of 31).
@@
Table 33: Statistics for domain: News Media (Part 21 of 31).
@@
Table 35: Statistics for domain: Other Manufacturing (Part 23 of 31).
@@
Table 41: Statistics for domain: Tourism Geography (Part 29 of 31).
@@
Table 42: Statistics for domain: Transportation (Part 30 of 31).
@@
Table 43: Statistics for domain: Water Resources Ocean (Part 31 of 31).

### v2

@@ 6 Conclusion
In this paper, we introduce PosIR, the first standardized benchmark for diagnosing position bias in diverse retrieval scenarios, covering 10 languages and 31 domains. Our experiments reveal prevalent primacy and unexpected recency biases across both multilingual and cross-lingual settings, while also exposing significant discrepancies with existing short-text benchmarks. These findings contribute new empirical evidence on position bias in dense retrieval, paving the way for more position-robust retrieval systems.
@@
Despite our efforts, this study has several limitations: 1) Although we employ LLMs to extract reference spans and implement a strict contrastive verification pipeline, the localization of relevant information is inherently approximate. LLM-extracted spans may not perfectly capture the minimal necessary evidence, and our contrastive filtering validates span necessity only indirectly—by confirming that removing the span degrades relevance scores. However, span removal may also disrupt the document’s overall semantic coherence, potentially conflating span necessity with document fluency degradation. As a result, a marginal risk remains that certain spans are imprecise or that alternative relevant spans are overlooked. 2) While LLM-based translation significantly extends the linguistic diversity of PosIR, it is intrinsically difficult to completely eliminate subtle semantic drift or translation artifacts across all datasets. Moreover, positional annotations derived from the source language may shift to varying degrees due to differences in linguistic structure (e.g., word order, morphological expansion), potentially introducing noise into position bias evaluation for translated languages. Although our sampling-based automatic and human evaluations indicate high fidelity, translation quality in certain low-resource languages might still be influenced by the underlying model bias. 3) Our empirical analysis focuses on embedding-based dense retrieval models. Although prior research [33] indicates that cross-encoders may exhibit superior position robustness, a systematic large-scale investigation of alternative architectures, including generative retrieval, remains for future work. Nevertheless, PosIR provides the necessary model-agnostic framework to facilitate such benchmarking.
Appendix for PosIR
For detailed information of the models appearing in this paper, please refer to Table 12.
Appendix 0.A PosIR Datasets
0.A.1 Overview
@@
PosIR comprises 310 datasets (17,329,673 documents and 421,708 queries) spanning 10 languages and 31 domains, forming a large-scale multilingual benchmark for information retrieval. For each dataset, we use the same format as BEIR, i.e. corpus, queries and qrels, which will be all available in the Hugging Face Hub. For each dataset, we use the same format as BEIR, i.e. corpus, queries and qrels, which are all available in the Hugging Face Hub. [^12] https://huggingface.co/datasets/infgrad/PosIR-Benchmark-v1 Table 4 summarizes language-level statistics, showing well-balanced corpus sizes across languages, with each language contributing approximately 1.7M–1.9M documents. Token length statistics reveal substantial cross-lingual variation that reflects differences in writing systems and morphological complexity. Logographic languages exhibit more compact representations (e.g., Chinese and English), whereas morphologically rich alphabetic languages require longer sequences (e.g., Arabic and Russian). Table 5 reports domain-level aggregates and highlights pronounced variation in domain scale and document characteristics. Domain sizes range from several hundred thousand documents to over three quarters of a million, while token length distributions reflect domain-specific content properties. Technical domains such as Information Security tend to contain longer documents, whereas more formal or symbolic domains, such as Mathematics and Statistics, exhibit comparatively shorter texts. This diversity enables evaluation of IR systems under heterogeneous conditions. For comprehensive coverage, Tables 1343 provide detailed statistics for all 31 domains, presented individually for each domain. These tables report language-specific document counts, query counts, average token lengths, and token length ratios relative to English. Such fine-grained reporting supports reproducibility and facilitates the analysis of domain–language interaction effects in multilingual IR.
Table 4: Summary of corpus statistics aggregated by language.
Document Count Token Length
Language Corpus Queries Ratio Min Q1 Median Q3 Max Mean
Arabic 1,712,847 41,894 9.89% 5 717 1401 2250 4719 1509.5
Chinese 1,870,228 43,931 10.78% 13 473 869 1410 2050 950.8
German 1,718,556 41,983 9.92% 3 655 1278 2067 4296 1387.3
English 1,719,701 42,015 9.92% 14 416 842 1392 2051 922.4
French 1,718,832 42,001 9.92% 2 645 1269 2054 4352 1375.1
Italian 1,718,831 42,005 9.92% 3 660 1295 2090 4298 1401.3
Korean 1,715,299 41,913 9.90% 5 704 1403 2258 4685 1509.5
Portuguese 1,718,848 42,002 9.92% 3 607 1195 1941 4020 1298.1
Russian 1,718,056 41,969 9.91% 2 721 1424 2300 4828 1539.2
Spanish 1,718,475 41,995 9.92% 4 604 1187 1925 3956 1287.9
Total/Avg 17,329,673 421,708 100.00% 5 620 1216 1968 3925 1318.1
Table 5: Summary of corpus statistics aggregated by domain.
@@
0.A.2 Diversity Analysis
0.A.2.1 Query Diversity
@@
0.A.2.2 Corpus Diversity
Following the approach in MTEB [18], we compute inter-domain similarities separately for English and Chinese corpora, each consisting of 31 domains, in PosIR. Specifically, we adopt the highest-scoring model on the MTEB (eng, v2) STS task, QZhou-Embedding (91.65) [31], to generate embeddings for all documents within each domain. We then average document embeddings at the domain level and compute pairwise cosine similarities between domains. The resulting similarity matrix is visualized as a heatmap in Figure 6.
@@ 0.A.2.2 Corpus Diversity
Appendix 0.B Human Evaluation of Generated Queries and Reference Spans
0.B.1 Evaluation Setup
To validate the quality of the generated queries and reference spans in PosIR, we conducted targeted human evaluation on a representative subset. Given the large scale of PosIR (310 datasets spanning 10 languages and 31 domains), we focused our human evaluation on the English subset to enable a focused and thorough assessment. Specifically, we randomly sampled 500 English query-document-reference span triples across 10 diverse domains (50 samples per domain). Each sample consists of: (1) a query generated by LLMs, (2) the corresponding positive document, and (3) the reference span annotated in the qrels, which we evaluated to verify whether it accurately localizes the answer within the document.
0.B.2 Evaluation Protocol
One annotator evaluated each sample across three dimensions using a structured evaluation form:
  • Query Quality (1-5 Scale): This dimension assesses the overall quality of the query for retrieval. Scores range from 5 (a perfect query that is clear, specific, fully relevant to the document) to 1 (query with significant issues in relevance, clarity, or answerability).
  • Span Necessity (Binary 0/1): This dimension evaluates whether the reference span contains only the necessary content. A score 1 indicates the span captures the core answer without irrelevant context, and a score 0 indicates over-extraction with unnecessary surrounding content.
  • Span Sufficiency (Binary 0/1): This dimension evaluates whether the reference span contains complete information to answer the query. A score 1 indicates the span is sufficient with the core answer completely contained, and a score 0 indicates under-extraction where some supporting details are missing.
We also evaluated whether the generated spans fabricate information not grounded in the document.
0.B.3 Evaluation Results
The evaluation results demonstrate the high quality of PosIR’s position-aware annotations. Query quality achieved a mean score of 4.91, with 496 of 500 samples (99.2%) rated $\geq$ 4, indicating Good or Excellent quality. For span precision, 490 of 500 spans (98.0%) satisfied both necessity and sufficiency criteria. Among the 10 remaining cases, 3 exhibited over-extraction and 7 exhibited under-extraction. Notably, no hallucinations were observed. These results confirm that PosIR’s fine-grained position-aware relevance annotations are reliable and suitable for rigorous evaluation of position bias in retrieval models [^14] We acknowledge that a single annotator limits the assessment of inter-rater reliability; however, the near-ceiling scores across all dimensions suggest that annotation quality is unambiguous in the vast majority of cases..
Appendix 0.C Multilingual Machine Translation
@@
We evaluate four representative translation systems: (1) Google Translate [^15] https://translate.google.com, a widely used commercial neural machine translation service; (2) Hunyuan-MT-7B [39], a specialized translation model achieved first place in the WMT25 competition; (3) Qwen3-30B-A3B-Instruct-2507 [30], a 31B-parameter open-source general MoE model; and (4) GPT-4o [20], a frontier general model serving as a strong baseline. Each system translates the sampled texts into 9 target languages: Arabic (ara-Arab), French (fra-Latn), German (deu-Latn), Italian (ita-Latn), Korean (kor-Kore), Portuguese (por-Latn), Russian (rus-Cyrl), and Spanish (spa-Latn). The prompt for the translation task is provided in Appendix 0.D.4.
0.C.1 Automatic Evaluation
0.C.1.1 Evaluation Protocol
To efficiently evaluate the translation quality across 8 languages and 31 domains, we employ an LLM-based reference-free evaluation approach [14] (see Appendix 0.D.5). Specifically, we utilize GPT-4o as an automatic evaluator to assess translation quality along four dimensions: Fluency (naturalness and readability of the target text), Accuracy (semantic fidelity to the source), Completeness (preservation of all source information), and Style Consistency (appropriateness of tone and register). Each dimension is rated on a 5-point Likert scale (1 = Poor, 5 = Excellent), and an overall score is computed as the mean of the four dimensions. In total, this results in 2,976 evaluated translation outputs (372 samples $\times$ 8 languages).
@@ 0.C.1.2 Evaluation Results
0.C.2 Human Evaluation
Initial inspection using automatic evaluation suggests that Qwen3-30B-A3B-Instruct-2507 produces high-quality translations across PosIR, achieving competitive performance in comparison to other methods. However, given that translation quality appears strong, rigorous human validation is essential. While automatic evaluation can detect catastrophic failures, it may be insensitive to subtle but critical errors in high-quality translations, such as nuanced mistranslations, contextual inappropriateness, or domain-specific inaccuracies. Due to limited resources, we conduct focused human evaluation on three typologically diverse languages (Arabic, French, and Russian) to provide expert assessment of translation quality, facilitating a direct comparison between human evaluation and LLM-based assessments.
@@ 0.C.2 Human Evaluation
0.C.2.1 Annotator Recruitment
Five annotators were recruited from graduate programs in linguistics, natural language processing, and applied foreign languages via established academic networks. All annotators were advanced graduate students or recent PhD graduates with expertise in computational linguistics and/or translation studies. For Arabic and Russian, native speakers of the respective languages conducted the annotations. For French, two annotators holding DALF C1/C2 certificates (Diplôme Approfondi de Langue Française, the highest level of French language certification) completed the evaluation. All annotators were compensated at a fair hourly rate, in compliance with applicable labor regulations.
0.C.2.2 Evaluation Protocol
@@
Annotators evaluated 372 translation pairs across 3 languages (1,116 samples in total) generated by Qwen3-30B-A3B-Instruct-2507. The evaluations were conducted in accordance with the consistency standards outlined in the automatic evaluation, as described in Appendix 0.C.1.1. Annotators were given both the English source text and the target language translation, with no access to reference translations to prevent anchoring bias.
0.C.2.3 Evaluation Results
@@
Table 9: Human-GPT-4o agreement within ±1 point tolerance.
@@
0.C.3 GPT-4o Human Evaluation Alignment
@@
0.C.4 Model Selection Rationale
@@
0.C.5 Fixing and Filtering Strategies
@@ 0.C.5.2 Refusal Detection
0.C.5.3 Statistical Outlier Filtering
We perform token-length-based outlier filtering at the language–domain level using the Qwen3 tokenizer. For each stratum, we compute the interquartile range (IQR) and define asymmetric thresholds, with a lower bound of $Q1-3.0\times IQR$ and an upper bound of $Q3+1.5\times IQR$. This asymmetric design focuses on filtering abnormally long texts, which are more likely to result from translation errors, duplication, or unintended document concatenation. It enforces a stricter upper bound to remove such cases, while allowing reasonable length variation among shorter documents. We process only corpus outliers, cascading deletions to associated queries and qrels.
In total, 15,641 corpus documents (0.11%) and 332 queries (0.10%) are removed. The largest reductions occur in Arabic (5,490 corpus entries; 0.25%) and Korean (4,082; 0.24%), followed by Russian (1,557; 0.11%). Overall, the removal ratios remain consistently low across all languages, indicating that the filtering procedure effectively eliminates anomalous long texts while preserving the structural integrity and cross-lingual balance of the benchmark.
Appendix 0.D Prompts
0.D.1 Prompt for Position-Aware QA Generation
@@
0.D.2 Prompt for Reference Locating
@@
0.D.3 Prompt for Relevance Assessment
Previous work [28] demonstrate that large language models such as GPT-4 can achieve labeling quality comparable to human annotators when producing gold-standard relevance labels for search systems. Moreover, [32] shows that incorporating fine-grained relevance labels into prompts for LLM-based rerankers significantly improves zero-shot reranking performance by better distinguishing borderline and ambiguous relevance cases. Motivated by these findings, we employ DeepSeek-V3.1 (Non-thinking Mode) [^18] https://api-docs.deepseek.com/news/news250821 as a relevance labeler, adopting a 5-level relevance labeling scheme to capture fine-grained distinctions in relevance. This labeling process is used solely for auxiliary verification rather than primary relevance annotation.
@@ 0.D.3 Prompt for Relevance Assessment
0.D.4 Prompt for Multilingual Machine Translation
@@
0.D.5 Prompt for Translation Quality Evaluation
@@
Qwen3 tokenizer [30] - https://huggingface.co/Qwen/Qwen3-30B-A3B-Instruct-2507/blob/main/tokenizer.json
@@
gte-multilingual-base [35] 305M https://huggingface.co/Alibaba-NLP/gte-multilingual-base
bge-m3 [3] 568M https://huggingface.co/BAAI/bge-m3
Qwen3-Embedding-0.6B [37] 595M https://huggingface.co/Qwen/Qwen3-Embedding-0.6B
inf-retriever-v1-1.5b 2B https://huggingface.co/infly/inf-retriever-v1-1.5b
Qwen3-Embedding-4B [37] 4B https://huggingface.co/Qwen/Qwen3-Embedding-4B
inf-retriever-v1 7B https://huggingface.co/infly/inf-retriever-v1
QZhou-Embedding [31] 7B https://huggingface.co/Kingsoft-LLM/QZhou-Embedding
NV-Embed-v2 [15] 8B https://huggingface.co/nvidia/NV-Embed-v2
llama-embed-nemotron-8b [1] 8B https://huggingface.co/nvidia/llama-embed-nemotron-8b
Qwen3-Embedding-8B [37] 8B https://huggingface.co/Qwen/Qwen3-Embedding-8B
KaLM-Embedding-Gemma3-12B-2511 [38] 12B https://huggingface.co/tencent/KaLM-Embedding-Gemma3-12B-2511
@@
bge-reranker-v2-m3 [3] 0.6B https://huggingface.co/BAAI/bge-reranker-v2-m3
Qwen3-Reranker-0.6B [37] 0.6B https://huggingface.co/Qwen/Qwen3-Reranker-0.6B
bge-reranker-v2-minicpm-layerwise [3] 3B https://huggingface.co/BAAI/bge-reranker-v2-minicpm-layerwise
Qwen3-Reranker-4B [37] 4B https://huggingface.co/Qwen/Qwen3-Reranker-4B
@@
Hunyuan-MT-7B [39] 8B https://huggingface.co/tencent/Hunyuan-MT-7B
Qwen3-30B-A3B-Instruct-2507 [30] 31B https://huggingface.co/Qwen/Qwen3-30B-A3B-Instruct-2507
DeepSeek-V3.1 [7] 685B https://huggingface.co/deepseek-ai/DeepSeek-V3.1
GPT-4o [20] - https://platform.openai.com/docs/models/gpt-4o
@@
Table 18: Statistics for domain: Biomedicine (Part 6 of 31).
@@
Table 22: Statistics for domain: Electric Power Energy (Part 10 of 31).
@@
Table 23: Statistics for domain: Film Entertainment (Part 11 of 31).
@@
Table 24: Statistics for domain: Finance Economics (Part 12 of 31).
@@
Table 25: Statistics for domain: Fineweb (Part 13 of 31).
@@
Table 28: Statistics for domain: Law Judiciary (Part 16 of 31).
@@
Table 29: Statistics for domain: Literature Emotion (Part 17 of 31).
@@
Table 33: Statistics for domain: News Media (Part 21 of 31).
@@
Table 35: Statistics for domain: Other Manufacturing (Part 23 of 31).
@@
Table 36: Statistics for domain: Petrochemical (Part 24 of 31).
@@
Table 41: Statistics for domain: Tourism Geography (Part 29 of 31).
@@
Table 42: Statistics for domain: Transportation (Part 30 of 31).
@@
Table 43: Statistics for domain: Water Resources Ocean (Part 31 of 31).